On January 1, 2017 PT Megaworld had 50,000 shares, $4 par value per share, during the year, the following transactions occurred. Apr Issued 9,000 additional shares of common stock for $11 per share. June 15 Declared a cash dividend of S1,50 per share to stockholders of record on June 30 July 10 Paid the $1.50 cash dividend. December Issued 4,000 additional shares of common stock for $12 per share, December 15 Declared a cash dividend on outstanding shares of $1.60 per share to stockholders of record on December 31 Required: Record all transactions above.
Answer #1
Debit Credit
Apr. 1 Cash 99000 =9000*11
      Common Stock 36000 =9000*4
      Paid in capital in excess of par-Common Stock 63000
June 15 Cash Dividends 88500 =(50000+9000)*1.5
       Dividends payable 88500
July 10 Dividends payable 88500
        Cash 88500
Dec 1 Cash 48000 =4000*12
      Common Stock 16000 =4000*4
      Paid in capital in excess of par-Common Stock 32000
Dec 15 Cash Dividends 100800 =(50000+9000+4000)*1.6
       Dividends payable 100800